satellite73 (satellite73):

@amistre64 please refresh my feeble memory snap way to find polynomial through (-2,4),(0,-6),(4,70)

OpenStudy (amistre64):

hmm, my method is to construct it such that each x value causes the unknown constants to zero out

OpenStudy (anonymous):

i thought i had this, but i resorted to a two by two system then i tried writing \[4-5(x+2)+4x(x+2)\] which worked but it was agony finding those constants i though you had a snap way of finding them

OpenStudy (amistre64):

if you can do a matrix augment, thats pretty snappy

OpenStudy (anonymous):

yeah but i wanted the snappy method you used

hero (hero):

....also interested in "snappy" method :D

OpenStudy (amistre64):

i believe the method you posted is what i used at first (-2,4),(0,-6),(4,70) y = a + bx + cx(x+2) ; (0,-6) -6 = a +0+0 y = -6 + bx + cx(x+2) ; (-2,4) 10 = -2b + 0 ; b=-5 y = -6 -5x + cx(x+2) ; (4,70) 70+6+20 = c4(4+2) 96/24 = c = 4 y = -6 -5x + 4x(x+2)

OpenStudy (anonymous):

Expanding this should work as well

OpenStudy (amistre64):

if there was another method that was snappier; you might have to refresh me memory how it looked to you

OpenStudy (amistre64):

joes looks kinda legendre to me

hero (hero):

@joemath314159, your method looks like some numerical methods stuff

OpenStudy (anonymous):

no that was what i was looking for i guess that is what i did, somehow it looked easier when you did it. i guess the grass is always greener hello @Hero hello @joemath314159

OpenStudy (anonymous):

It is legendre, learned it in Linear Algebra

OpenStudy (amistre64):

its these lighntning fast clickers of mine, just makes it appear rico y suave lol

OpenStudy (anonymous):

(0,-2) (1,1) (2,6) (3,19)

hero (hero):

@satellite73, are you taking a course?

OpenStudy (anonymous):

yes, at the school of amistre

hero (hero):

lol

OpenStudy (anonymous):

\[a+bx+cx(x-1)+cx(x-1)(x-2)+dx(x-1)(x-2)\] actually i think this is slightly different then legrange, but maybe it is identical

OpenStudy (amistre64):

(0,-2) (1,1) (2,6) (3,19) y = -2 + bx +cx(x-1) + dx(x-1)(x-2) 1 = -2 + b; b= 3 y = -2 + 3x +cx(x-1) + dx(x-1)(x-2) 6 = -2 +6 + 2c ; c=1 y = -2 + 3x +x(x-1) + dx(x-1)(x-2) 19 = -2 + 9 + 6 + d(2); d=3 y = -2 + 3x +x(x-1) + 3x(x-1)(x-2) if i mathed it right

OpenStudy (amistre64):

legendre i believe zeros out each constant term and omits one zero so that just a single constant is exposed at any time. Newtons method is similar to mine in that new information can be added as needed without have to reconstruct the whole equation

OpenStudy (anonymous):

yeah legendre looks like what @joemath314159 wrote newtons i am not sure about

OpenStudy (anonymous):

now i can sleep better. thanks!